# Continuous fields of interval algebras

### v1

@@ 4. An outro on continuous fields of Cuntz semigroups
The final theorem and corollary of the previous section classify a specific class continuous fields of $C^{*}$-algebras and its morphisms. It has been common to express such results in a categorical fashion, via a well-defined functor $F\colon\mathcal{C}\rightarrow\mathcal{D}$. (See e.g. [6, Section 2].) This section suggests a categorical context $\mathcal{F}_{\operatorname{Cu}}$ of continuous fields of Cuntz semigroups, suitable to construct a well-defined classifying functor.
@@
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ be a continuous field of $C^{*}$-algebras. Observe that $\operatorname{Cu}(A_{x})$ and $\operatorname{Cu}(\mathcal{A})$ are $\operatorname{Cu}$-semigroups and that for any $x\in X$, we have a surjective $\operatorname{Cu}$-morphism $\operatorname{Cu}(\pi_{x})\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(A_{x})$. Also, There is a natural order-embedding $\nu\colon\operatorname{Cu}(\mathcal{A})\lhook\joinrel\rightarrow\operatorname{Cu}-\prod_{x}\operatorname{Cu}(A_{x})$ given by $[a]\mapsto\overline{p_{a}}$ where $p_{a}$ is the $\ll_{\rm{pw}}$-increasing path of $\prod_{x}\operatorname{Cu}(A_{x})$ given by $\epsilon\mapsto(\operatorname{Cu}(\pi_{x})([(a-\epsilon)_{+}]))_{x}$. Next, we deduce (iii) of Section 4 from [1, Lemma 2.2 (i)], while we left (iv) for the reader to check.
@@
A fortiori, any $\mathcal{F}_{\operatorname{Cu}}$-isomorphism $\alpha\colon\operatorname{Cu}((X,\{A_{x}\}_{x},\mathcal{A}))\rightarrow\operatorname{Cu}((X,\{B_{x}\}_{x},\mathcal{B}))$ lift to a $\mathcal{F}_{C^{*}}$-isomorphism $\phi\colon(X,\{A_{x}\}_{x},\mathcal{A})\rightarrow(X,\{B_{x}\}_{x},\mathcal{B})$ which is unique up to approximate unitary equivalence.

### v2

@@ 4. An outro on continuous fields of Cuntz semigroups
The final theorem and corollary of the previous section classify a specific class continuous fields of $C^{*}$-algebras and its morphisms. It has been common to express such results in a categorical fashion, via a well-defined functor $F\colon\mathcal{C}\rightarrow\mathcal{D}$. (See e.g. [11, Section 2].) This section suggests a categorical context $\mathcal{F}_{\operatorname{Cu}}$ of continuous fields of Cuntz semigroups, suitable to construct a well-defined classifying functor.
@@
Let $(X,\{A_{x}\}_{x},\mathcal{A})$ be a continuous field of $C^{*}$-algebras. Observe that $\operatorname{Cu}(A_{x})$ and $\operatorname{Cu}(\mathcal{A})$ are $\operatorname{Cu}$-semigroups and that for any $x\in X$, we have a surjective $\operatorname{Cu}$-morphism $\operatorname{Cu}(\pi_{x})\colon\operatorname{Cu}(\mathcal{A})\rightarrow\operatorname{Cu}(A_{x})$. Also, There is a natural order-embedding $\nu\colon\operatorname{Cu}(\mathcal{A})\lhook\joinrel\rightarrow\operatorname{Cu}-\prod_{x}\operatorname{Cu}(A_{x})$ given by $a]\mapsto\overline{p_{a}}$ where $p_{a}$ is the $\ll_{\rm{pw}}$-increasing path of $\prod_{x}\operatorname{Cu}(A_{x})$ given by $\epsilon\mapsto(\operatorname{Cu}(\pi_{x})([(a-\epsilon)_{+}]))_{x}$. Next, we deduce (iii) of Section 4 from [1, Lemma 2.2 (i)], while we left (iv) for the reader to check.
@@
A fortiori, any $\mathcal{F}_{\operatorname{Cu}}$-isomorphism $\alpha\colon\operatorname{Cu}((X,\{A_{x}\}_{x},\mathcal{A}))\rightarrow\operatorname{Cu}((X,\{B_{x}\}_{x},\mathcal{B}))$ lifts to a unique (up to approximate unitary equivalence) $\mathcal{F}_{C^{*}}$-isomorphism $\phi\colon(X,\{A_{x}\}_{x},\mathcal{A})\rightarrow(X,\{B_{x}\}_{x},\mathcal{B})$.